# Large Language Models Are Zero-Shot Time Series Forecasters

### v1

@@ 5 Origins of Zero-Shot Performance
To understand why LLMs can extrapolate time series in a zero-shot manner, let’s take a step back and consider simple numerical sequences, for example [$1,4,9,16,\dots$] or [$0,0,1,0,1,2,0,1,2,3,\dots$]. For any input sequence, there are arbitrarily many generation rules that are consistent with the input (e.g. $f(x)=x^{2}$ for $x\in[1,2,3,4,...]$), but some generation rules are overly complex and will generalize poorly. LLMs can forecast effectively because they prefer completions derived from simple rules, adopting a form of Occam’s razor prior (Goldblum et al., 2023; Sutskever, 2023; Delétang et al., 2023). To explicitly demonstrate this phenomenon, we create a synthetic example using the function $f(x)=x+\cos(x)$ with additive Gaussian noise. We fit symbolic expressions to the first 70% of timesteps using PySR (Cranmer, 2023) with symbols ["$+$", $\cdot$", "-", "///", "$\sin$", "$\cos$", "$\exp$","square"] to identify generating rules with known complexity, quantified by the number of symbols in the regressed expression (Appendix D). Figure 5 (left) shows the likelihood that GPT-3 assigns the highest likelihood to symbolic regression generating rules that balance consistency with complexity.
@@
  • Repetition bias and periodicity: LLMs’ bias towards repetitive sequences (Holtzman et al., 2019) (often unwanted in NLP) corresponds precisely to the ability to identify and extrapolate periodic structure in the input. $4.2,8.6,1.0,4.2,8.6$ will lead to a $1.0$ as a likely next output without any time series or arithmetic knowledge ($x_{t}=x_{t-T})$.
  • Arithmetic and trend components: LLMs’ ability to perform addition and multiplication (Yuan et al., 2023; Liu and Low, 2023) maps on to extrapolating linear and exponential trends. For example, predicting the next element of 0.2,1.6,3.,4.40.2,1.6,3.,4.40.2 , 1.6 , 3 . , 4.4 the LLM needs only to add $1.4$ to the last element ($x_{t+1}=x_{t}+c$). Similarly, exponential trends have the generating rule $x_{t+1}=c\cdot x_{t}$ and sigmoid trends have the generating rule $x_{t+1}=x_{t}+cx_{t}(1-x_{t})$.
Combining multiple patterns together presents a more difficult challenge, as it requires both identifying the composite pattern and being able to perform the multiple operations within the same token budget. Supposing that a model can perform copying in a single forward pass and addition in a single forward pass, that does not necessarily mean that it can do both simultaneously. We find that GPT-3 is only sometimes able to perform these compositions, though GPT-4 does so more consistently as shown in Appendix E. It is likely that the limitations on compute and tokens spent may make this composition unnecessarily hard, and that additional recursive structure, for example from a scratchpad (Nye et al., 2021), Chain of Thought (CoT) prompting (Wei et al., 2022), or adaptive computation (Schwarzschild et al., 2021; Anil et al., 2022), would make this task easier.
figureLLMs can find low complexity explanations of the data, enabling them to zero-shot extrapolate numerical sequences. (left) GPT-3 likelihoods favor solutions from symbolic regression (PySR (Cranmer, 2023)) that balance training loss and complexity, leading to good generalization. (right) GPT-3 predicted median and 10-90th percentile prediction interval are shown given 140 timesteps of context. On the right of each time series, we show the log likelihoods compared to the ARIMA and TCN time series models. Overall, GPT-3 performs considerably better than the baselines, though composition and exponential growth are more challenging for the models (Appendix D.1).
@@ 6 Special Properties of LLMs
Base models and forecasting performance    Given the rapid growth and improvement in open-source LLMs (Touvron et al., 2023a, b), the relationship between LLMTime forecasting performance and the performance of the underlying base model is particular important and has broad implications. Steady increases in LLM benchmark performance can directly translate to steady improvements in forecasting ability. In Figure 6 (right), we show a study with OpenAI models (davinci, babbage, curie, ada), variants of LLaMA (Touvron et al., 2023a) (7B, 13B, 33B, 65B) and LLaMA-2 (Touvron et al., 2023b) models (7B, 13B, 70B) measuring accuracy on the Massive Multitask Language Understanding (MMLU) benchmark and probabilistic forecasting error. As we might hope, when reasoning (MMLU) performance increases forecasts also improve.
@@
Chat models    Though convenient scaling relationships appear to hold for base models, the correlation begins to break down when we consider models that have been post-processed for chatbot applications. GPT-4 (OpenAI, 2023), for example, demonstrates considerably more intelligence than GPT-3 and LLaMA models in natural language tasks, but effectively applying it to time series is challenging. In Figure 6 (center), we show that GPT-4 has a forecasting error (CRPS) significantly larger than GPT-3’s on the Darts datasets. The performance drop is the result of several small details in GPT-4’s method. Due to the altered tokenization, GPT-4 cannot be easily forced to tokenize individual digits into an unbroken stream of numbers. Due to the restrictions on the API, likelihood evaluations are also not permitted, which is why we present results for only CRPS. While GPT-4 can perform well on the synthetic examples discussed in Section 5 (shown in Appendix E), we find that the GPT-4 calibration is much worse than GPT-3’s on stochastic data, likely as a result of the preprocessing details above and the fact that the model has been treated with RLHF (Christiano et al., 2017) which is known to degrade calibration on question-answering tasks (OpenAI, 2023). GPT-4 is not the only example of degraded performance in models designed for chat functionality. We observed the same phenomenon in LLaMA-2 models, which have corresponding chat versions for every model size. Figure 6 (right) shows that chat versions tend to have markedly worse forecasting error than their non-chat counterparts, though still maintain trends in size and reasoning ability.
@@
Figure 6: (left) Time series forecasting performance (NLL/D and CRPS on Darts (Herzen et al., 2022)) improves with reasoning performance of the underlying model LLM, as judged by accuracy on the Massive Multitask Language Understanding (MMLU) benchmark (Hendrycks et al., 2020). Displayed results are for all GPT-3, LLaMA (Touvron et al., 2023a), and LLaMA-2 (Touvron et al., 2023b) base models. (center) GPT-4 performs worse than GPT-3 (right) Forecasting performance (NLL/D and CRPS on Darts) appears to be negatively affected by alignment procedures (e.g. instruction tuning and RLHF) in general. LLaMA-2 chat models typically perform worse than the corresponding base model. Error bars show standard errors over individual datasets.
@@
Connecting time series and textual understanding    Because LLMs are designed for natural language and code, we can augment the numerical time series with useful text. We can do so either by providing textual side information as inputs, or by producing textual outputs from a given time series. An interesting question is whether GPT-4 can explain in text its understanding of a given time series. We probe this quality by providing GPT-4 the code to generate our synthetic time series, provide the values of one these time series, and then ask it to infer which of the functions produced the data in a zero-shot manner. The prediction accuracies are shown in Figure 7, with the three remaining rows all being 00. With CoT (Wei et al., 2022) prompting the model performs much better than random chance; however, its ability to identify patterns better when directly extrapolating the numerical data, suggesting that its numerical understanding is not fully connected to its textual understanding. In making predictions, the model often explains properties of the time series in order to select the right candidate from the list, and we show several of these sample explanations in Appendix F. We also show how this task is encapsulated in a simple (unprompted) next token prediction problem on cells of a Jupyter notebook, illustrating why we expect such capabilities to emerge with a sufficiently powerful language model.
Figure 7: (left) LLMTime can handle missing values without interpolation by denoting missingness with text (e.g. ‘NaN’). For baseline methods we perform linear interpolation and then fit the model as usual. LLMTime assigns higher log likelihood to datasets preprocessed with added ‘NaN’s than baseline methods assign to interpolated datasets. Forecasting performance, as judged by CRPS, is competitive between LLMTime and alternative methods that use explicit interpolation. Filled area shows standard error over individual datasets and 3 random seeds. (right) LLMs can be used to answer questions about time series data posed as text. We show GPT-4’s accuracy at predicting the function that generated the time series, obtained using chain-of-thought prompting.
@@ 7 Discussion
We have demonstrated that large language models can be used as pretrained time series forecasters by encoding numerical values as text. As with other “foundation” models, pretraining confers useful biases toward generalizable patterns that would otherwise be engineered into the model through architecture design (Gruver et al., 2022), and enables natural scaling of performance with improvements in the base pretrained model. Because LLM forecasters are trained on language, they also confer unconventional capabilities, such as question answering. More broadly, framing time series forecasting as natural language generation can be seen as another step towards unifying more capabilities within a single large and powerful model, in which understanding can be shared between many tasks and modalities. Moreover, zero-shot forecasting can enable broadly compelling performance without requiring significant computational resources, domain expertise, or many downstream training data points.
While LLM forecasters benefit from the strengths of pretrained transformers, they also inherit their weaknesses, which can include a limited context window. While many univariate time series problems can fit comfortably within increasingly large context windows, multivariate problems pose a more significant challenge. There have been several recent advances extending LLM context windows to 10-100K tokens (OpenAI, 2023; Anil et al., 2023; Anthropic, 2023; Ainslie et al., 2023). Combining these advances with time series forecasting is a particularly exciting direction for future research. Another potential challenge of using current LLMs architectures could be their weakness in arithmetic and performing recursive and compositional operations, which could be a limitation on particularly challenging time series. On the other hand, many time series do not require precise arithmetic. Understanding the extent to which this is the case, and relaxing this limitation, is also a promising avenue for future research. Separately from any limitation, it would also be promising to investigate effective procedures for fine-tuning LLMs on time series. We hope that bridging LLM research with time series forecasting brings benefits to both communities.
Acknowledgements. We thank Micah Goldblum, Greg Benton, and Wesley Maddox for helpful discussions. This work is supported by NSF CAREER IIS-2145492, NSF I-DISRE 193471, NIH R01DA048764-01A1, NSF IIS-1910266, NSF 1922658 NRT-HDR, Meta Core Data Science, Google AI Research, BigHat Biosciences, Capital One, and an Amazon Research Award.
@@ Appendix B Addressing Memorization Concerns in GPT-3 Evaluations
Evaluating the performance of black box APIs, like those provided by OpenAI, can be challenging when training data for the underlying models is unknown. In our time series setting, it is natural to wonder if the common benchmark datasets we use are included in the GPT-3 training data. LLMs are known to memorize large amounts of their training data verbatim, including common benchmark text datasets and copyrighted material [Biderman et al., 2023, Chang et al., 2023]. Beyond outright memorization, more benign data leakage of closely related data is also possible, leading to overestimation of the generalization performance.
Even if our evaluation datasets are present in the GPT-3 training data, it’s unlikely that GPT-3’s good performance is the result of memorization for at least two reasons a priori. First of all, our idiosyncratic formatting is unlikely to be present in the training dataset, even if the numerical values and their order are the same. Second, the time series datasets are unlikely to appear in GPT-3’s training data sufficiently frequently to lead to memorization, as memorization increases in proportion with redundancy [Lee et al., 2021].
@@ C.1 Darts datasets
For the Darts datasets, we use the GPyTorch library [Gardner et al., 2018] for Gaussian Process implementation and the DARTS libary [Herzen et al., 2022] for ARIMA, TCN, N-BEATS, N-HiTS. We use default values for hyperparameters not described below. The test set is the last $20\%$ of each series.
We use several baseline methods implemented directly in Darts [Herzen et al., 2022]:
  • ARIMA: ARIMA [Box and Jenkins, 1968], short for AutoRegressive Integrated Moving Average, has been a popular choice for time series forecasting for many decades.
  • TCN: Temporal Convolutional Network (TCN) [Lea et al., 2016] is residual network with dilated 1D convolutions.
  • N-BEATS: N-BEATS [Oreshkin et al., 2020] is a deep learning model tailored for time series forecasting. It employs a deep architecture with backward and forward residual links and stacked fully-connected layers.
  • N-HiTS: N-HiTS [Challu et al., 2022] is a deep learning model that incorporates hierarchical interpolation and multi-rate data sampling techniques in order to create forecasts that emphasize different frequencies and scales of the input signal.
We also include Spectral Mixture Gaussian Process (SM-GP) [Wilson and Adams, 2013] as a Bayesian nonparametric approach to time series modeling.

### v2

@@ 5 Origins of Zero-Shot Performance
To understand why LLMs can extrapolate time series in a zero-shot manner, let’s take a step back and consider simple numerical sequences, for example [$1,4,9,16,\dots$] or [$0,0,1,0,1,2,0,1,2,3,\dots$]. For any input sequence, there are arbitrarily many generation rules that are consistent with the input (e.g. $f(x)=x^{2}$ for $x\in[1,2,3,4,...]$), but some generation rules are overly complex and will generalize poorly. LLMs can forecast effectively because they prefer completions derived from simple rules, adopting a form of Occam’s razor prior [19, 42, 15]. To explicitly demonstrate this phenomenon, we create a synthetic example using the function $f(x)=x+\cos(x)$ with additive Gaussian noise. We fit symbolic expressions to the first 70% of timesteps using PySR [14] with symbols ["$+$", $\cdot$", "-", "///", "$\sin$", "$\cos$", "$\exp$","square"] to identify generating rules with known complexity, quantified by the number of symbols in the regressed expression (Appendix D). Figure 5 (left) shows the likelihood that GPT-3 assigns the highest likelihood to symbolic regression generating rules that balance consistency with complexity.
@@
  • Repetition bias and periodicity: LLMs’ bias towards repetitive sequences [25] (often unwanted in NLP) corresponds precisely to the ability to identify and extrapolate periodic structure in the input. $4.2,8.6,1.0,4.2,8.6$ will lead to a $1.0$ as a likely next output without any time series or arithmetic knowledge ($x_{t}=x_{t-T})$.
  • Arithmetic and trend components: LLMs’ ability to perform addition and multiplication [51, 30] maps on to extrapolating linear and exponential trends. For example, predicting the next element of 0.2,1.6,3.,4.40.2,1.6,3.,4.40.2 , 1.6 , 3 . , 4.4 the LLM needs only to add $1.4$ to the last element ($x_{t+1}=x_{t}+c$). Similarly, exponential trends have the generating rule $x_{t+1}=c\cdot x_{t}$ and sigmoid trends have the generating rule $x_{t+1}=x_{t}+cx_{t}(1-x_{t})$.
Combining multiple patterns together presents a more difficult challenge, as it requires both identifying the composite pattern and being able to perform the multiple operations within the same token budget. Supposing that a model can perform copying in a single forward pass and addition in a single forward pass, that does not necessarily mean that it can do both simultaneously. We find that GPT-3 is only sometimes able to perform these compositions, though GPT-4 does so more consistently as shown in Appendix E. It is likely that the limitations on compute and tokens spent may make this composition unnecessarily hard, and that additional recursive structure, for example from a scratchpad [33], Chain of Thought (CoT) prompting [47], or adaptive computation [41, 3], would make this task easier.
[…]
figureLLMs can find low complexity explanations of the data, enabling them to zero-shot extrapolate numerical sequences. Left: GPT-3 likelihoods favor solutions from symbolic regression (PySR [14]) that balance training loss and complexity, leading to good generalization. Right: GPT-3 predicted median and 10-90th percentile prediction interval are shown given 140 timesteps of context. On the right of each time series, we show the log likelihoods compared to the ARIMA and TCN time series models. Overall, GPT-3 performs considerably better than the baselines, though composition and exponential growth are more challenging for the models (Appendix D.1).
@@ 6 Special Properties of LLMs
Base models and forecasting performance    Given the rapid growth and improvement in open-source LLMs [43, 44], the relationship between LLMTime forecasting performance and the performance of the underlying base model is particular important and has broad implications. Steady increases in LLM benchmark performance can directly translate to steady improvements in forecasting ability. In Figure 6 (right), we show a study with OpenAI models (davinci, babbage, curie, ada), variants of LLaMA [43] (7B, 13B, 33B, 65B) and LLaMA-2 [44] models (7B, 13B, 70B) measuring accuracy on the Massive Multitask Language Understanding (MMLU) benchmark and probabilistic forecasting error. As we might hope, when reasoning (MMLU) performance increases forecasts also improve.
@@
Chat models    Though convenient scaling relationships appear to hold for base models, the correlation begins to break down when we consider models that have been post-processed for chatbot applications. GPT-4 [36], for example, demonstrates considerably more intelligence than GPT-3 and LLaMA models in natural language tasks, but effectively applying it to time series is challenging. In Figure 6 (center), we show that GPT-4 has a forecasting error (CRPS) significantly larger than GPT-3’s on the Darts datasets. The performance drop is the result of several small details in GPT-4’s method. Due to the altered tokenization, GPT-4 cannot be easily forced to tokenize individual digits into an unbroken stream of numbers. Due to the restrictions on the API, likelihood evaluations are also not permitted, which is why we present results for only CRPS. While GPT-4 can perform well on the synthetic examples discussed in Section 5 (shown in Appendix E), we find that the GPT-4 calibration is much worse than GPT-3’s on stochastic data, likely as a result of the preprocessing details above and the fact that the model has been treated with RLHF [13] which is known to degrade calibration on question-answering tasks [36]. GPT-4 is not the only example of degraded performance in models designed for chat functionality. We observed the same phenomenon in LLaMA-2 models, which have corresponding chat versions for every model size. Figure 6 (right) shows that chat versions tend to have markedly worse forecasting error than their non-chat counterparts, though still maintain trends in size and reasoning ability.
@@
Figure 6: Left: Time series forecasting performance (NLL/D and CRPS on Darts [23]) improves with reasoning performance of the underlying model LLM, as judged by accuracy on the Massive Multitask Language Understanding (MMLU) benchmark [22]. Displayed results are for all GPT-3, LLaMA [43], and LLaMA-2 [44] base models. Center: GPT-4 performs worse than GPT-3. Right: Forecasting performance (NLL/D and CRPS on Darts) appears to be negatively affected by alignment procedures (e.g. instruction tuning and RLHF) in general. LLaMA-2 chat models typically perform worse than the corresponding base model. Error bars show standard errors over individual datasets.
Missing data    A major advantage of treating forecasting as a text completion task and using LLMs is that we can easily feed in any input that can be encoded as text. Often in time series, the time series will be incomplete and certain observations are missing. Simple imputation methods, such as nearest neighbor, are still core pre-processing steps in common data science workflows [32], and the choice of imputation method is especially relevant to clinical data, which often contains irregularly sampled measurements and where missingness can be meaningful signal in itself [27]. Much like humans reading partial reports, LLMs can handle missing values without imputation by adopting special symbols, for instance,
@@
. In Figure 7 we compare likelihoods and CRPS value for forecasts from traditional time series methods and LLaMA-2 70B on data that has been corrupted with missing values and then processed with linear interpolation and the above string formatting. While the likelihoods of traditional methods rapidly deteriorate with corruptions, we find that LLaMA-2 70B is more resilient, and when comparing CRPS values, LLaMA-2 70B is competitive with methods that use interpolation.
Connecting time series and textual understanding    Because LLMs are designed for natural language and code, we can augment the numerical time series with useful text. We can do so either by providing textual side information as inputs, or by producing textual outputs from a given time series. An interesting question is whether GPT-4 can explain in text its understanding of a given time series. We probe this quality by providing GPT-4 the code to generate our synthetic time series, provide the values of one these time series, and then ask it to infer which of the functions produced the data in a zero-shot manner. The prediction accuracies are shown in Figure 7, with the three remaining rows all being 00. With CoT [47] prompting the model performs much better than random chance; however, its ability to identify patterns better when directly extrapolating the numerical data, suggesting that its numerical understanding is not fully connected to its textual understanding. In making predictions, the model often explains properties of the time series in order to select the right candidate from the list, and we show several of these sample explanations in Appendix F. We also show how this task is encapsulated in a simple (unprompted) next token prediction problem on cells of a Jupyter notebook, illustrating why we expect such capabilities to emerge with a sufficiently powerful language model.
Figure 7: Left: LLMTime can handle missing values without interpolation by denoting missingness with text (e.g. ‘NaN’). For baseline methods we perform linear interpolation and then fit the model as usual. LLMTime assigns higher log likelihood to datasets preprocessed with added ‘NaN’s than baseline methods assign to interpolated datasets. Forecasting performance, as judged by CRPS, is competitive between LLMTime and alternative methods that use explicit interpolation. Filled area shows standard error over individual datasets and 3 random seeds. Right: LLMs can be used to answer questions about time series data posed as text. We show GPT-4’s accuracy at predicting the function that generated the time series, obtained using chain-of-thought prompting.
@@ 7 Discussion
We have demonstrated that large language models can be used as pretrained time series forecasters by encoding numerical values as text. As with other “foundation” models, pretraining confers useful biases toward generalizable patterns that would otherwise be engineered into the model through architecture design [21], and enables natural scaling of performance with improvements in the base pretrained model. Because LLM forecasters are trained on language, they also confer unconventional capabilities, such as question answering. More broadly, framing time series forecasting as natural language generation can be seen as another step towards unifying more capabilities within a single large and powerful model, in which understanding can be shared between many tasks and modalities. Moreover, zero-shot forecasting can enable broadly compelling performance without requiring significant computational resources, domain expertise, or many downstream training data points.
While LLM forecasters benefit from the strengths of pretrained transformers, they also inherit their weaknesses, which can include a limited context window. While many univariate time series problems can fit comfortably within increasingly large context windows, multivariate problems pose a more significant challenge. There have been several recent advances extending LLM context windows to 10-100K tokens [36, 4, 5, 1]. Combining these advances with time series forecasting is a particularly exciting direction for future research. Another potential challenge of using current LLMs architectures could be their weakness in arithmetic and performing recursive and compositional operations, which could be a limitation on particularly challenging time series. On the other hand, many time series do not require precise arithmetic. Understanding the extent to which this is the case, and relaxing this limitation, is also a promising avenue for future research. Separately from any limitation, it would also be promising to investigate effective procedures for fine-tuning LLMs on time series. We hope that bridging LLM research with time series forecasting brings benefits to both communities.
Acknowledgements. We thank Micah Goldblum, Greg Benton, and Wesley Maddox for helpful discussions. This work is supported by NSF CAREER IIS-2145492, NSF I-DISRE 193471, NSF IIS-1910266, BigHat Biosciences, Capital One, and an Amazon Research Award.
@@ A.2 Validation tuning
A.3 Likelihood adjustment for GPT Models
In order to convert token probabilities assigned by language models into continuous densities, we must convert the distribution over all possible tokens into the distribution over only tokens used in the numerical encoding scheme. When we have access to the language model’s logits, performing this adjustment is easy. We can simply set the probability of any non-essential tokens to zero and renormalize the distribution. When using black-box APIs (e.g. the OpenAI API), however, we need to approximate this procedure, becauase it frequently impossible to obtain the full discrete distribution over tokens. For example, in the OpenAI API, only the top 5 log probabilities are returned for every step in the generation process. As we cannot properly renormalize the distribution, we have to make adjustments that get us as close as possible to the true renormalized distribution, as we show in the following calculations. We take $p$ to be the raw probabilities assigned by a language model and $\tilde{p}$ to be the adjusted probabilities, with
$p_{k}=\frac{e^{\log p_{k}}}{\sum_{i}e^{\log p_{i}}}\,\,,\hskip 14.22636pt\tilde{p}_{k}=\frac{e^{\log p_{k}}}{\left(\sum_{i}e^{\log p_{i}}\right)-e^{l_{0}}}$
where $l_{0}$ is the probability that mass assigned to tokens that are not part of the numerical encoding scheme. When access to log probabilities is limited, $l_{0}$ can be approximated as the sum of the log probabilities for non-numerical tokens in the top $k$. From the definition of $\tilde{p}$ we can derive
$\frac{1}{\tilde{p}}$ $=e^{l_{0}}+\frac{1}{p}\implies$
$\tilde{p}$ $=\left(\frac{1}{p}-e^{l_{0}}\right)^{-1}\implies$
$\log\tilde{p}$ $=\log p-\log(1-e^{l_{0}})$
@@ Appendix B Addressing Memorization Concerns in GPT-3 Evaluations
Evaluating the performance of black box APIs, like those provided by OpenAI, can be challenging when training data for the underlying models is unknown. In our time series setting, it is natural to wonder if the common benchmark datasets we use are included in the GPT-3 training data. LLMs are known to memorize large amounts of their training data verbatim, including common benchmark text datasets and copyrighted material [7, 12]. Beyond outright memorization, more benign data leakage of closely related data is also possible, leading to overestimation of the generalization performance.
Even if our evaluation datasets are present in the GPT-3 training data, it’s unlikely that GPT-3’s good performance is the result of memorization for at least two reasons a priori. First of all, our idiosyncratic formatting is unlikely to be present in the training dataset, even if the numerical values and their order are the same. Second, the time series datasets are unlikely to appear in GPT-3’s training data sufficiently frequently to lead to memorization, as memorization increases in proportion with redundancy [29].
@@ C.1 Darts datasets
For the Darts datasets, we use the GPyTorch library [17] for Gaussian Process implementation and the DARTS libary [23] for ARIMA, TCN, N-BEATS, N-HiTS. We use default values for hyperparameters not described below. The test set is the last $20\%$ of each series.
We use several baseline methods implemented directly in Darts [23]:
  • ARIMA: ARIMA [8], short for AutoRegressive Integrated Moving Average, has been a popular choice for time series forecasting for many decades.
  • TCN: Temporal Convolutional Network (TCN) [28] is residual network with dilated 1D convolutions.
  • N-BEATS: N-BEATS [37] is a deep learning model tailored for time series forecasting. It employs a deep architecture with backward and forward residual links and stacked fully-connected layers.
  • N-HiTS: N-HiTS [11] is a deep learning model that incorporates hierarchical interpolation and multi-rate data sampling techniques in order to create forecasts that emphasize different frequencies and scales of the input signal.
We also include Spectral Mixture Gaussian Process (SM-GP) [48] as a Bayesian nonparametric approach to time series modeling.